Estimate: 0.5162 + 9.436 + 350.12 + 800.25 + 25.78 =

3004 x 5982 =

Answer by checkley77(12569)

0.5162 + 9.436 + 350.12 + 800.25 + 25.78 =1+9+350+800=1,160 ans.
3004 x 5982 =3000*6000=18,000,000 ans.